OpenStudy (anonymous):

Given tGiven the following probability distribution, find P (X = 3). X 1 2 3 4 5 P(X) 0.4 0.1 0.15 0.2 (1 point) 2. The number of cars an American family owns follows the distribution below: Number of Cars 0 1 2 3 4 5 Probability 0.09 0.36 0.35 0.13 0.05 0.02 a) Verify that this is a legitimate probability distribution. b) Interpret (in words) the notation P(X > 2). c) Interpret (in words) the notation P(X = 2). d) Find P(X > 2). (5 points) 3. A study of social mobility in America examined the social class attained by the sons of lower class fathers. Social classes were numbered from 1 to 5 with 1 representing the lower class and 5 the higher class. Consider the random variable X to the class of a randomly chosen son. The study found the following distribution: Son's Class 1 2 3 4 5 Probability 0.48 0.38 0.08 0.05 0.01 a) What percent of the sons reached the highest class? b) Check that this distribution meets the requirements of a discrete probability distribution. c) What is P(X < 2)? d) What is P(X = 2)? e) Write the event: a son of a lower-class father attains one of the highest two social classes in terms of X. (6 points)

ganeshie8 (ganeshie8):

a trick for getting quick help : limit to one question per post

ganeshie8 (ganeshie8):

for 1st question, use below : sum of probabilities must add up to 1

ganeshie8 (ganeshie8):

X 1 2 3 4 5 P(X) 0.4 0.1 0.15 0.2 <<<<<<<<<

ganeshie8 (ganeshie8):

that row must add up to "1"

ganeshie8 (ganeshie8):

0.4 + 0.1 + P(3) + 0.15 + 0.2 = 1

ganeshie8 (ganeshie8):

solve P(3)

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

np :) did u solve P(3) ha...

OpenStudy (anonymous):

so add all the numbers together?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

0.4 + 0.1 + P(3) + 0.15 + 0.2 = 1 P(3) + 0.85 = 1 P(3) = 0.15

ganeshie8 (ganeshie8):

see if that makes more or less sense :)

OpenStudy (anonymous):

15

OpenStudy (anonymous):

0.15

ganeshie8 (ganeshie8):

*its 0.15

ganeshie8 (ganeshie8):

yup

ganeshie8 (ganeshie8):

hey i need to go... please post the second question in a separate post... good luck !

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

np :)